# Parameter-Efficient Fine-Tuning with Attributed Patch Semantic Graph for Automated Patch Correctness Assessment

### v1

@@ 4.1 Overview
Motivation: Previous works have demonstrated the importance of capturing patch semantic and explicitly considering certain code attributes in statically predicting patch correctness [62, 38, 73, 33, 64]. While the proposed patch graph representation APSG adequately captures such information, we further need to effectively incorporate information of APSG into LLMs for statically predicting patch correctness. Inspired by the work of Yao et al. [72], we find that the attention mechanism can effectively merge the graph information in APSG with the sequence information in LLMs. Besides, LLMs need to be fine-tuned in order to adapt to the APCA task. Taking these aspects into account, on top of LoRA [28]—one of the most advanced LLM parameter-efficient fine-tuning (PEFT) methods, we propose a new PEFT method called Graph-LoRA to retain graph information and fully train LLMs. Graph-LoRA can effectively fine-tune the parameters of LLMs and incorporate APSG information into LLMs through the attention mechanism.
Figure 2: An overview of fine-tuning LLM with Graph-LoRA.
Framework: Fig. 2 shows the process of fine-tuning LLMs with Graph-LoRA. Given the buggy line(s), patched line(s) and context of the buggy code, the specific process of our method is as follows: (a) First, we preprocess the patch data to obtain the APSG and sequence representation of the patch; (b) Then, we obtain the graph features and sequence features of the patch through GNN and LLM respectively; (c) Next, we use the attention mechanism of Graph-LoRA to merge the graph features with sequence features; (d) Finally, LLM processes the merged features and assesses the correctness of the patch.
@@ 4.2 Code Embedding for Sequence Features
LLMs can convert the patch into token embeddings that will be used for prediction in subsequent modules. In this work, we use LLMs built by the stacked decoder of transformers [58], which is the most popular LLM in the field of software engineering.
Given an input sequence $X$ of a code piece containing the patch, let $X_{i}$ be the i-th token of the code piece. To make the LLMs clearly distinguish the patch content, we use pre-appended token $<P>$ to wrap the beginning and end of the patch. In addition, we add the text ”Assess whether the patch is correct” in the front of the code piece, which serves as a prompt to LLM. Finally, the code piece with patch is represented as:
@@ 4.3 GNN for Graph Features
To enrich the feature vector of the patch, we need to additionally get the feature of the APSG. According to the procedure in Section 3, we can build the APSG of the patch and extract the node matrix $N$, adjacency matrix $M$, and attribute matrix $A$. The node matrix includes the line node matrix $N_{l}$ and the variable node matrix $N_{v}$, and the attribute matrix includes the attributes of each node. The adjacency matrix includes the line node adjacency matrix $M_{l}$, the variable node adjacency matrix $M_{v}$, and the sub-graph merge edge matrix $M_{l\_v}$. The line node matrix, line node adjacency matrix, and line node attribute matrix form the overall graph. The variable node matrix, variable node adjacency matrix, and variable node attribute matrix form the subgraph. To effectively obtain the graph information of APSG, given the powerful ability of the Graph Neural Network (GNN) [48], we make use of GNN to process graph data and extract features of APSG. The process of extracting feature of APSG is as follows: (a) First, we process node features and node attribute features; (b) Then, based on the node and its attribute features, we extract subgraph features; (c) Finally, we pass the subgraph features to the overall graph and extract the overall graph features.
@@ 4.4 Graph-LoRA
First, Graph-LoRA decomposes the weight update matrix $\Delta W$ of LLM into low-order matrixs $W_{down}$ and $W_{up}$ through low-rank decomposition. The weights of the LLM are updated as follows:
$\Delta W=W_{down}W_{up}$ (12)
Second, based on the APSG features generated by GNN, Graph-LoRA merges the graph information in APSG with the sequence information in LLMs through the attention mechanism. Graph-LoRA aims to use the information in APSG to help LLMs assess the correctness of patches. The specific operations are as follows:
@@
$\Delta W=W_{down}Attention(E,F_{APSG})W_{up}$ (14)
where $GNN$ represents the graph neural network, $F_{APSG}$ represents the APSG features of patch, $E$ represents the sequence features of patch. Specifically, the attention mechanism mainly consists of three weight matrices Q, K, and V. The matrix Q introduces external information to obtain the attention score by calculating the dot product between it and the matrix K, and the obtained attention score can distinguish the importance of external information. We find that the external information introduced by the matrix Q can guide the original features to change. Based on this idea, we use the matrix Q in the multi-head attention mechanism to introduce the graph information of APSG and guide the update of patch features in the LLM. The detailed operations are as follows:
$\text{Attention}(\Delta W,F_{APSG})=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W_{O}$ (15)
$\text{head}_{i}(\Delta W,F_{APSG})=S(\Delta W,F_{APSG})(\Delta WV)$ (16)
$S(\Delta W,F_{APSG})=\text{softmax}\left(\frac{(F_{APSG}Q)(\Delta WK^{T})}{\sqrt{d_{k}}}\right)$ (17)
where $Q$, $K$, $V$, $W_{O}$ are weight matrices. During fine-tuning, the initial weight $W_{0}$ of LLM is frozen and does not receive gradient updates, but the parameters in $W_{down}$,$W_{up}$, the attention fusion layer, and GNN are updated. During backpropagation, LLM obtains the information of APSG by adding the new parameter update matrix $\Delta W$ as follows:
$W=W_{0}+\Delta W$ (18)
@@ 4.5 Training and Inference
During training, the parameters in Graph-LoRA and GNN are trained jointly. The additional training parameters of Graph-LoRA are equivalent to 0.6% of the LLM parameters of 7B size, thus keeping the training cost low. Following the previous studies [82, 73, 38], we perform standard 10-fold cross-validation during training. We use the cross entropy loss to calculate the gap between the model results and the true value, which has been widely used in classification tasks and previous APCA task. We continuously reduce the gap between the true label y and the model prediction result p to update the model parameters. The cross-entropy loss operation is as follows:
@@ 5.1 Research Questions
RQ1 (Effectiveness): How does our model perform compared to other existing works on the APCA task? To pursue this question, we evaluate the model on four widely used datasets in the APCA task and compare the performance with that of the state-of-the-art APCA methods.
@@ 5.2 Datasets
In this study, we focus on APCA task datasets constructed with patches for Defect4J subjects for the following reasons: (1) Defect4J contains hundreds of real-world bugs for widely used, large-scale real-world projects, and is the most widely used APR dataset for studying APR techniques in the literature; (2) most existing APCA works are evaluated using Defect4J, so our experimental results can be more fairly compared with previous work. More specifically, we select four Defect4J-related APCA task datasets, and Table II gives a summary of these datasets. These datasets range from big to medium to small in terms of size, rang from balanced to imbalanced in terms of the ratio between the correct and overfitting patches. We next give a brief description of these four selected datasets.
@@
Datasets # Correct # Overfitting Total
Wang 248 654 902
Merge 271 2,489 2,760
Balance 271 271 542
Lin 535 648 1,183
@@
Wang dataset. The Wang dataset is the most widely used dataset for the APCA task. Wang et al. [60] use 21 mainstream APR tools to fix bugs in Defects4J V1.2 and collect the patches generated by these tools. They then check the collected patches and manually assess their correctness. For the patches that pass the tests, they mark them either as correct or overfitting. Finally, they obtain a total of 902 patches, including 248 correct patches and 654 overfitting patches.
Lin dataset. Compared with the Wang dataset, the Lin dataset contains more patches. To better explore the overfitting problem, Lin et al. [38] add patches generated by some other well-known APR tools such as JAID, SketchFix, CapGen, SOFix, and SequenceR to the Wang dataset. To avoid data leakage, they remove duplicate patches, ultimately obtaining 1183 patches.
@@
Merge dataset. The Merge dataset is the largest manually labeled dataset for the APCA task. Yang et al. [70] manually label the 1,988 patches generated by the PraPR repair system [20] and merge them with the Wang dataset by carefully removing the duplicates. They finally obtain 2,760 patches, including 2,489 overfitting patches and 271 correct patches.
@@
Balance dataset. The Balance dataset contains an equal number of overfitting patches and correct patches. For more thorough evaluations, Yang et al. [70] construct the Balance dataset to address the problem that the number of correct patches is too different from that of the overfitting patches in the Merge dataset. More specifically, they keep all correct patches from the Merge dataset and sample the same number of overfitting patches from the Merge dataset.
@@
Overall, the four selected datasets include the largest manually labeled dataset—the Merge dataset, the relatively small but most widely used dataset—the Wang dataset, the imbalanced datasets—the Wang dataset and the Merge dataset, and the (relatively) balanced datasets—the Lin dataset and the Balance dataset. Collectively, using these four datasets enable us to conduct a thorough and comprehensive evaluation.
Note while there exist datasets provided by Ye et al. [73] and Lin et al. [38] that contain around 10,000 and 50,000 patches respectively, these patches are not manually labeled and using them can potentially threat the validity of the experimental results as pointed out by Yang et al. [70]. To reduce the threat, we thus select the four datasets which contain only accurately labeled patches.
@@ 5.3 Baselines
Currently, APCA methods are mainly divided into dynamic methods and static methods. We select representative works in the two categories as our baselines.
@@
Among the dynamic methods, we select two representative works as our baselines: PATCH-SIM [67] and Opad [71]. PATCH-SIM exploits the behavior similarity of test case executions, and is currently the best among dynamic methods. Opad uses fuzz testing to generate new test cases for exposing overfitting patches, and Opad is further divided into E-Opad and R-Opad according to the different test generation tools used (Evosuite vs Randoop). To facilitate fair comparison, we refer to the results by Yang et al.[70] about PATCH-SIM and Opad for relevant datasets.
@@
Static methods can be further divided into three categories: traditional methods (denoted as static-traditional), machine learning based methods (denoted as static-ML), and large language model based methods (denoted as static-LLM). For traditional methods, we consider S3 [33], ssFix [61], and CapGen [62]. For machine learning based methods, we consider ODS [73], CACHE [38], and APPT [82]. Moreover, we consider a large language model based method LLM4PatchCorrect [83]. Among traditional methods, S3 assesses the correctness of patches based on six features, ssFix assesses the patch correctness based on structural and conceptual information, and finally CapGen assesses the patches based on contextual information. Among the machine learning-based methods, ODS extracts 202 patch features through abstract syntax trees to describe the correct patch, CACHAE considers both the changed code segments and the related unchanged code segments, APPT adopts a pre-trained model as an encoder stack and then uses an LSTM stack and a deep learning classifier to evaluate patch correctness. LLM4PatchCorrect predicts the patch correctness by feeding a LLM with information of labeled patches, such as error descriptions and failed tests.
Similar to dynamic methods, we refer to results by Yang et al.[70] about S3, ssFix, CapGen, and ODS for relevant datasets. For CACHE and APPT, their original papers report the results for the Lin dataset and we refer to these results for fair comparison, and we strictly follow the corresponding artifacts to run them and get their results for the other three considered datasets. Previous works have shown that machine learning-based methods significantly outperform traditional methods in APCA tasks, and APPT is the most advanced APCA method based on machine learning. More specifically, according to the experimental results of Zhang et al. [82], the performance of APPT is better than ODS, CACHAE, and other machine learning based methods like BATS [52]. LLM4PatchCorrect is the most advanced APCA method based on LLM [83], and it also outperforms all other traditional methods and machine learning based methods. To give a more extensive evaluation, we account for three representative LLMs in this paper: CodeLlama, StarCoder, and Llama3. As the original LLM4PatchCorrect paper [83] just reports the result obtained using the StarCoder LLM, we strictly follow the methodology described in the paper to obtain the results for the other two LLMs. We also implement our approach using these three LLMs, and evaluate our approach for the considered datasets.

### v2

@@ 4 Graph-LoRA for LLMs
Figure 2: An overview of fine-tuning LLM with Graph-LoRA in our method.
@@ 4.1 Overview
Motivation: Previous works have demonstrated the importance of capturing patch semantic and explicitly considering certain code attributes in statically predicting patch correctness [wen2018context, lin2022context, ye2021automated, le2017s3, xia2023automated]. While the proposed patch graph representation APSG adequately captures such information, we further need to effectively incorporate information of APSG into LLMs for statically predicting patch correctness. Inspired by the work of Yao et al. [yao2020multimodal], we find that the attention mechanism can effectively merge the graph information in APSG with the sequence information in LLMs. Besides, LLMs need to be fine-tuned in order to adapt to the APCA task. Taking these aspects into account, on top of LoRA [hu2021lora]—one of the most advanced LLM parameter-efficient fine-tuning (PEFT) methods, we propose a new PEFT method called Graph-LoRA to retain graph information and fully train LLMs. Graph-LoRA can effectively fine-tune the parameters of LLMs and incorporate APSG information into LLMs through the attention mechanism.
Framework: Fig. 2 shows the process of fine-tuning LLMs with Graph-LoRA in our method. Given the buggy line(s), patched line(s) and context of the buggy code, the specific process of our method is as follows: (a) First, we preprocess the patch data to obtain the APSG and sequence representation of the patch; (b) Then, we obtain the graph features and sequence features of the patch through GNN and LLM respectively; (c) Next, we use the attention mechanism of Graph-LoRA to merge the graph features with sequence features; (d) Finally, LLM processes the merged features and assesses the correctness of the patch.
@@ 4.2 Code Embedding for Sequence Features
LLMs can convert the patch into token embeddings that will be used for prediction in subsequent modules. In this work, we use LLMs built by the stacked decoder of transformers [vaswani2017attention], which is the most popular LLM in the field of software engineering.
Given an input sequence $X$ of a code piece containing the patch, let $X_{i}$ be the i-th token of the code piece. To make the LLMs clearly distinguish the patch content, we use pre-appended token <P><P> to wrap the beginning and end of the patch. In addition, we add the text ”You are a model responsible for assessing patch correctness. Assess whether the patch is correct” in the front of the code piece, which serves as a prompt to LLM. Finally, the code piece with patch is represented as:
@@ 4.3 GNN for Graph Features
To enrich the feature vector of the patch, we need to additionally get the feature of the APSG. According to the procedure in Section 3, we can build the APSG of the patch and extract the node matrix $N$, adjacency matrix $M$, and attribute matrix $A$. The node matrix includes the line node matrix $N_{l}$ and the variable node matrix $N_{v}$, and the attribute matrix includes the attributes of each node. The adjacency matrix includes the line node adjacency matrix $M_{l}$, the variable node adjacency matrix $M_{v}$, and the sub-graph merge edge matrix $M_{l\_v}$. The line node matrix, line node adjacency matrix, and line node attribute matrix form the overall graph. The variable node matrix, variable node adjacency matrix, and variable node attribute matrix form the subgraph. To effectively obtain the graph information of APSG, given the powerful ability of the Graph Neural Network (GNN) [scarselli2008graph], we make use of GNN to process graph data and extract features of APSG. The process of extracting feature of APSG is as follows: (a) First, we process node features and node attribute features; (b) Then, based on the node and its attribute features, we extract subgraph features; (c) Finally, we pass the subgraph features to the overall graph and extract the overall graph features. We next give details of the three steps.
@@ 4.4 Graph-LoRA
First, Graph-LoRA decomposes the pre-trained weights $W_{0}\in R^{d\times k}$ of LLMs into a vector $m\in R^{1\times k}$ representing “magnitude” and a matrix $V\in R^{d\times k}$ representing “direction” on top of DoRA [liu2024dora], an advanced variant of LoRA. Compared to the original LoRA method that fine-tunes all content in one step, this method allows LLMs to clearly know the magnitude and direction of the weights that need to be fine-tuned, resulting in better performance. For the magnitude vector, we fine-tune all its parameters using the weight update vector $\Delta m$. For the direction matrix, we decompose its update matrix $\Delta V$ into low-rank matrices $V_{down}\in R^{d\times r}$ and $V_{up}\in R^{r\times k}$ using low-rank decomposition. The specific operation is as follows:
$m=\left\lVert W_{0}\right\rVert_{c}\quad\quad V=W_{0}$ (12)
$\Delta V=V_{down}V_{up}$ (13)
where $\lVert.\rVert_{c}$ is the vector-wise norm of a matrix across each column vector.
Second, the low-rank matrices $V_{up}$ and $V_{down}$ compress the original weight dimension from $k$ to $r$, which may cause important information in the features to be lost. Therefore, Graph-LoRA uses PiSSA [meng2024pissa] to initialize the two low-rank matrices $V_{up}$ and $V_{down}$ in order for preserving the most important $r$ features. Specifically, PiSSA performs singular value decomposition on the original $V$ and takes the first $r$ principal singular components. The first $r$ components represent the $r$ most important features. The low-rank matrices $V_{up}$ and $V_{down}$ are initialized to the same subspace based on the singular value decomposition results, so that their product is initially along the dominant direction of the LLM features. The specific operation is as follows:
$V=USX^{T}$ (14)
$V_{down}=U_{[:,\,:r]}\,S^{1/2}_{[:r,\,:r]}\quad\quad V_{up}=S^{1/2}_{[:r,\,:r]}\,X^{\top}_{[:,\,:r]}$ (15)
where $U\in R^{d\times min(d,k)}$ and $X\in R^{k\times min(d,k)}$ are the singular vectors with orthonormal columns, $S\in R^{min(d,k)\times min(d,k)}$ is a diagonal matrix with the singular values arranged in descending order on the diagonal, and $X^{T}$ is the transpose of $X$. The matrix slicing notations are the same as those in PyTorch, and $[:r]$ and $[:]$ denote the first r dimensions and all dimensions respectively.
Third, based on the APSG features generated by GNN, Graph-LoRA merges the graph information in APSG with the sequence information in LLMs through the attention mechanism. Graph-LoRA aims to use the information in APSG to help LLMs assess the correctness of patches. The specific operations are as follows:
@@
$\Delta V=V_{down}Attention(E,F_{APSG})V_{up}$ (17)
where $GNN$ represents the graph neural network, $F_{APSG}$ represents the APSG features of patch, and $E$ represents the sequence features of patch. Specifically, the attention mechanism use weight matrix $W_{Q}$ to map external information (APSG features of patch, $F_{APSG}$) to query vectors and use weight matrices $W_{K}$ and $W_{V}$ to map internal information (sequence features of patch, $E$) to key vectors and value vectors respectively. The attention scores are computed by taking the dot product between the query vectors and key vectors, followed by a softmax that normalizes these attention scores into a probability distribution. The computed attention scores highlight the most relevant positions and are then used to form a weighted sum of the value vectors, so that the model focuses on the most relevant external information. Yao et al. [yao2020multimodal] found that attention can guide the original features to acquire external information. Inspired by this work, we use the $W_{Q}$ in the multi-head attention mechanism to introduce the graph information of APSG and guide the update of patch features in the LLM. The detailed operations are as follows:
$\text{Attention}(E,F_{APSG})=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W_{O}$ (18)
$\text{head}_{i}(E,F_{APSG})=S(E,F_{APSG})(EW_{V})$ (19)
$S(E,F_{APSG})=\text{softmax}\left(\frac{(F_{APSG}W_{Q})(EW_{K}^{T})}{\sqrt{d_{k}}}\right)$ (20)
where $W_{O}$ is a weight matrix. During fine-tuning, the parameters of LLMs are updated via $\Delta m$ and $\Delta V$. The specific operations for updating LLM parameters are as follows:
$W^{\prime}=(m+\Delta m)\,\frac{V+\Delta V}{\left\lVert V+\Delta V\right\rVert_{c}}$ (21)
@@ 4.5 Training and Inference
During training, the parameters in Graph-LoRA and GNN are trained jointly. The additional training parameters of Graph-LoRA are equivalent to 0.6% of the LLM parameters of 7B size, thus keeping the training cost low. Following the previous studies [zhang2024appt, ye2021automated, lin2022context], we perform 10-fold cross-validation and take the average of 10 rounds of each training and testing process as the final performance of our method. We use the cross-entropy loss to calculate the gap between the model results and the true value, which has been widely used in classification tasks and previous APCA task. We continuously reduce the gap between the true label y and the model prediction result p to update the model parameters. The cross-entropy loss operation is as follows:
@@ 5.1 Research Questions
RQ1 (Effectiveness): How does our model perform compared to other existing works on the APCA task? To pursue this question, we evaluate the model on five APCA datasets and compare the performance with that of the state-of-the-art APCA methods.
@@ 5.2 Datasets
In this study, we focus on APCA task datasets constructed with patches for real-world projects.
@@
Datasets Benchmarks # Correct # Overfitting Total
Wang Defects4J V1.2 248 654 902
Merge Defects4J V2.0 271 2,489 2,760
Balance Defects4J V2.0 271 271 542
Lin Defects4J V2.0 535 648 1,183
Multi-
Benchmarks
Defects4J V2.0,
Bus.jar, Bears
2673 5121 7794
More specifically, we select five APCA task datasets and Table II gives a summary of these datasets. These datasets range in size from small to large, vary from one to multiple in terms of the number of bug benchmarks used for constructing the datasets, and vary from balanced to imbalanced in terms of the ratio between the number of correct patches and the number of overfitting patches. Next, we give a brief description of these five datasets.
@@
Wang dataset. The Wang dataset is the most widely used dataset for the APCA task. Wang et al. [wang2020automated] use 21 mainstream APR tools to fix bugs in Defects4J V1.2 and collect the patches generated by these tools. They then check the collected patches and manually assess their correctness. For the patches that pass the tests, they mark them either as correct or overfitting. Finally, they obtain a total of 902 patches, including 248 correct patches and 654 overfitting patches.
@@
Merge dataset. The Merge dataset is the largest manually labeled dataset for the APCA task. Yang et al. [yang2023large] manually label the 1,988 patches generated by the PraPR repair system [10.1145/3293882.3330559] and merge them with the Wang dataset by carefully removing the duplicates. They finally obtain 2,760 patches, including 2,489 overfitting patches and 271 correct patches.
@@
Balance dataset. The Balance dataset contains an equal number of overfitting patches and correct patches. For more thorough evaluations, Yang et al. [yang2023large] construct the Balance dataset to address the problem that the number of correct patches is too different from that of the overfitting patches in the Merge dataset. More specifically, they keep all correct patches from the Merge dataset and sample the same number of overfitting patches from the Merge dataset.
Lin dataset. Compared with the Wang dataset, the Lin dataset contains more patches. To better explore the overfitting problem, Lin et al. [lin2022context] add 1,000 patches released by Tian et al. [tian2020evaluating] to the Wang dataset. These 1,000 patches include patches generated by well-known APR tools such as JAID, SketchFix, CapGen, SOFix, and SequenceR, as well as patches written by Defects4J developers. To avoid data leakage, they remove duplicate patches, ultimately obtaining 1183 patches.
Multi-Benchmarks dataset. Since the above four datasets are all constructed with patches for Defects4J, to enable more comprehensive evaluation, we construct a new large patch dataset that involves patches for three bug benchmarks: Defects4J [just2014defects4j], Bugs.jar [saha2018bugs], and Bears [madeiral2019bears]. The specific construction process is as follows. First, we add data from the Lin and Merge datasets (note that these two datasets contain all patches from the Wang and Balance datasets) to the Multi-Benchmarks dataset. Then, to avoid an excessive number of overfitting patches for Defects4J, we add human-written patches for Defects4J as the correct patches to the Multi-Benchmarks dataset. Next, to supplement the patches for bug benchmarks besides Defects4J, we augment the Multi-Benchmarks dataset with the patches for bug benchmarks Bugs.jar and Bears released by Ye et al. [ye2021automated]. The patches released by Ye et al. [ye2021automated] consist of human-written patches and patches generated by 11 APR tools. Finally, after deduplication, the Multi-Benchmarks dataset contains 7794 patches, including 2673 correct patches and 5121 overfitting patches. Overall, the patches in the Multi-Benchmarks dataset were both human-written and automatically generated by 22 different APR tools. Note that Lin et al. [lin2022context] and Ye et al. [ye2021automated] similarly use both human-written and automatically generated patches to evaluate APCA methods. Table III summarizes the Multi-Benchmarks dataset.
@@
Overall, the five selected datasets include the largest manually labeled dataset—the Merge dataset, the relatively small but most widely used dataset—the Wang dataset, the imbalanced datasets—the Wang dataset and the Merge dataset, the (relatively) balanced datasets—the Lin dataset and the Balance dataset, and the large dataset that involves multiple bug benchmarks—the Multi-Benchmarks dataset. Collectively, using these five datasets enables us to conduct a thorough and comprehensive evaluation.
TABLE III: Summary of the Multi-Benchmarks dataset.
Benchmark Subjects Correct Overfitting All
Defects4J Merge dataset [yang2023large] 271 2489 2760
Lin dataset [lin2022context] 535 648 1183
Human-written 798 0 798
Bugs.jar Ye et al. [ye2021automated] 986 2275 3261
Bears Ye et al. [ye2021automated] 219 531 750
Total 2809 5943 8752
Total (deduplicated) 2673 5121 7794
@@ 5.3 Baselines
Following existing studies [yang2023large, zhang2024appt, lin2022context], our study selects existing state-of-the-art techniques that are designed for or can be adapted to the APCA task. The selected techniques can be broadly categorized into two categories, including static and dynamic techniques. In particular, we obey the flowing two strategies widely used by existing works [yang2023large, zhang2024appt, lin2022context]. First, we only include techniques that do not rely on the ground-truth patches (i.e., the oracle information) since the ground-truth patch information is unavailable for real-world bug fixing [wang2020automated, overfittingsurvey, yang2023large]. Second, we only include techniques designed for Java language as Java is the most targeted language in the APR community and the existing patches of real-world bugs are usually available in Java language.
@@
Among the dynamic methods, we select two representative works as our baselines: PATCH-SIM [xiong2018identifying] and Opad [yang2017better]. PATCH-SIM exploits the behavior similarity of test case executions, and is currently the best among dynamic methods. Opad uses fuzz testing to generate new test cases for exposing overfitting patches, and Opad is further divided into E-Opad and R-Opad according to the different test generation tools used (Evosuite vs Randoop).
@@
Static methods can be further divided into three categories: traditional methods (denoted as static-traditional), machine learning based (ML-based) methods (denoted as static-ML), and large language model based (LLM-based) methods (denoted as static-LLM). For traditional methods, we consider S3 [le2017s3], ssFix [weimer2013leveraging], and CapGen [wen2018context]. For ML-based methods, we consider ODS [ye2021automated], BERT-LR [tian2020evaluating], BATS [tian2022predicting], PANTHER [tian2023best], CACHE [lin2022context], and APPT [zhang2024appt]. Moreover, we consider an LLM-based method LLM4PatchCorrect [zhou2024leveraging]. Among traditional methods, S3 assesses the correctness of patches based on six features, ssFix assesses the patch correctness based on structural and conceptual information, and finally CapGen assesses patches based on contextual information. Among the ML-based methods, ODS extracts 202 patch features through abstract syntax trees to describe the correct patch, CACHAE considers both the changed code segments and the related unchanged code segments, and APPT adopts a pre-trained model as an encoder stack and then uses an LSTM stack and a deep learning classifier to evaluate patch correctness. The other three ML-based methods are proposed by Tian and his co-authors. In particular, BERT-LR uses representation learning techniques to learn code change embeddings to assess patch correctness, BATS is an unsupervised method that predicts patch correctness by checking the behaviors of patches against the specifications of failing tests, and PANTHER investigates the advantages of learning code representation and evaluates the correctness of patches by integrating learned embeddings with engineered features. LLM4PatchCorrect is the state-of-the-art method based on LLMs, which predicts patch correctness by feeding an LLM with information of labeled patches, such as error descriptions and failed tests.
Regarding the results of baselines, we reuse the results of baselines from recently published works [zhang2024appt, yang2023large, ye2021automated, zhou2024leveraging] to facilitate a fair comparison whenever it is possible. For experimental completeness, we also reproduce several state-of-the-art APCA methods on the studied datasets that were not considered in previous works. Specifically, we refer to the results by Yang et al. [yang2023large] about dynamic methods PATCH-SIM, E-Opad, and R-Opad and traditional methods S3, ssFix, and CapGen (of the static method category) for relevant datasets. For ML-based method ODS, the work by Yang et al. [yang2023large] reports results for the Wang dataset and the work by Zhang et al. [zhang2024appt] reports results for the Lin dataset. For ML-based methods CACHE and APPT, the work by Zhang et al. [zhang2024appt] reports results for the Lin dataset. For LLM-based method LLM4PatchCorrect, the work by Zhou et al. [zhou2024leveraging] reports results for the Lin dataset using the LLM StarCoder. Likewise, we refer to these results for the purpose of fair comparison.
We next give how we choose APCA methods to reproduce and obtain their results on the studied datasets that were not considered in previous works. For dynamic methods, they have limited practical applications due to the disadvantage of requiring significant time to generate and/or execute tests [zhang2024appt, zhou2024leveraging]. Besides, Patch-SIM can only be applied to Defects4J v1.2 (note that besides the Wang dataset, the other 4 studied datasets all involve Defects4J v2.0) and both Patch-SIM and Opad have a low recall for overfitting patches [yang2023large, wang2020automated]. Thus, we do not reproduce these two dynamic methods. For static methods, previous works have shown that ML-based methods significantly outperform traditional methods [zhang2024appt] and LLM-based methods outperform traditional methods and ML-based methods [zhou2024leveraging]. In accordance with this, we consider BERT-RL, BATS, PANTHER, Cache, APPT, and LLM4PatchCorrect to be the most advanced APCA methods proposed recently. To ensure the comprehensiveness of our experimental results, we reproduce these advanced APCA methods on the studied datasets that were not considered in previous works. Specifically, since BERT-RL, BATS, and PANTHER have not been considered by previous works for the studied datasets, we strictly reproduce them according to the corresponding artifacts and obtain their performance on the five datasets. We also reproduce Cache, APPT, and StarCoder-based LLM4PatchCorrect and obtain results on the four studied datasets (excluding the Lin dataset) not considered by previous works.
To give a more extensive evaluation, we account for two other representative LLMs besides StarCoder (used by LLM4PatchCorrect in [zhou2024leveraging]) when a method involves LLMs: CodeLlama and Llama3. More specifically, we implement two variants of LLM4PatchCorrect (i.e., CodeLlama-based LLM4PatchCorrect and Llama3-based LLM4PatchCorrect) and evaluate them on the five studied datasets. We also implement our method using the three LLMs StarCoder, CodeLlama, Llama3, and conduct the evaluation on the five studied datasets.